OpenStudy (anonymous):

You want to order a rectangular bulletin board to mount on a classroom wall that has an area of 90sq ft. Fire code requirements allow for no more then 30% of a classroom be covered by said board. If the length is 3 times as long as the width, what are the dimensions of the largest board that meets the requirements?

OpenStudy (campbell_st):

let w = width 3w = length so the area of 90 sq ft is equal to 90 = w(3x) or 90 = 3w^2 you now need to solve for w

OpenStudy (campbell_st):

once you have the width then triple it to get the length

OpenStudy (anonymous):

area of the largest board =90*30/100=27. length=l,width=w ie,l=3w then 3w^2=27 this gives w=3 and l=9, hence its dmensions are 9*3 sqft.

OpenStudy (anonymous):

I think I understand... is there any tricks to breaking all this down?